OpenStudy (anonymous):

I need to solve for the missing variable for a triangle A=48in squared h= b=12in and if you could explain how to solve the problem that would be excellent.

OpenStudy (jdoe0001):

got picture?

OpenStudy (anonymous):

Find the missing variable for a triangle: A = 48 i n^2 h = b = 12 in 10 in 36 in 8 in 4 in

OpenStudy (solomonzelman):

You are given that the Area f the triangle is 48 inch². And the base is equal to 12 inch. \(\LARGE\color{blue}{ A~∆ = \frac{b \times h}{2} }\) then plug in your given values and solve for height. \(\LARGE\color{blue}{ 48 = \frac{12 \times h}{2} }\)

OpenStudy (solomonzelman):

can you finish solving for the h (height) ?

OpenStudy (jdoe0001):

so.... do you know what "A" stand for?

OpenStudy (jdoe0001):

I mean, there is a background to this I assume I don't think you were just handed out random material

OpenStudy (anonymous):

the A stands for Area

OpenStudy (jdoe0001):

well.. then check SolomonZelman's lines then